Show that the space L(R) of all bounded real sequence is a linear space over a vector field R


1
Expert's answer
2021-11-01T13:31:24-0400

Let "\overline{x}=(x_1,x_2,...),\space \overline{y}=(y_1,y_2,...)\in L(R)" be two elements from the set L(R), We define their sum as "\overline{x}+\overline{y}=(x_1+y_1,x_2+y_2,...,x_i+y_i,..),i\in N" or as sum for each component. We should ptove, that "\overline x+\overline {y}\in L(R)" but it is easy because "\forall i\in N |x_i+y_i|\le |x_i|+|y_i|\le ||\overline{x}||_{L(R)}+||\overline{y}||_{L(R)}\rarr\\
||\overline{x}+\overline{y}||_{L(R)}\le ||\overline{x}||_{L(R)}+||\overline{y}||_{L(R)}<\infty"

Axioms of addition

"\overline{x}+\overline{y}=\overline{y}+\overline{x};\\
\overline{x}+(\overline{y}+\overline{z})=(\overline{x}+\overline{y})+\overline{z}"

are valid because it is true in R for each component of vectors.

In linear space must be zero element and we define it by the formula

"\overline 0=(0,0,0,..)" or vector with zeros only.

Next property takes place

"\forall \overline{x} \in L(R)\space \overline{x}+\overline 0=\overline x"

because it is true for each component.

Operation of multiplication elements of L(R) by scalar from R is defined by the formula

"(c\cdot \overline x)_i=c\cdot x_i,i\in N,\space c\in R,\space \overline{x}\in L(R)"

We have "c\cdot \overline x\in L(R)" and "||c\cdot \overline {x}||_{L(R)}\le |c|\cdot ||\overline{x}||_{L(R)}"

Operation of multiplication is defined through components , therefore next properties take place:

"c\cdot (\overline x+\overline y)=c\cdot \overline{x}+c\cdot \overline{y};\\
c\cdot (d\cdot \overline x)=(c\cdot d)\cdot \overline {x},\space c, d\in R"

"1\cdot \overline(x)=\overline{x}"

The opposite to "\overline{x}" element "-\overline{x}" we define as "(-\overline{x})_i=-x_i,i\in N"

We have:

"(-\overline{x})+\overline {x}=\overline {0},\\
-\overline{x}=(-1)\cdot \overline{x}"

Thus all axioms of linear space for L(R) have verified by us.